Circular No. 9030

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2009 F2 (McNAUGHT) R. H. McNaught reports his discovery of another comet on CCD images

taken with the 0.5-m Uppsala Schmidt telescope at Siding Spring (discovery observation tabulated below), the object appearing well condensed with a 12′′ coma and a possible very weak 15′′ tail in p.a. 260^{o} (involved with stars). Follow-up observations by McNaught on Mar. 20.49 UT show a 9′′ coma and no tail. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, R. Ligustri (Udine, Italy) reports that seven co-added 180-s CCD exposures (taken remotely with a 0.35-m f/7 telescope at Grove Creek Observatory, Trunkey, N.S.W., Australia on Mar. 20.6) show a coma diameter of ∼ 18′′.

2009 UT α_{2000} δ_{2000} Mag. Observer
Mar. 19.57501 10^{h}57^{m}44.^{s}69 −43^{o}26′24′′.7 17.7 McNaught

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2009-F47.

T = 2009 Nov. 20.793 TT  ω = 338.^{o}758
Ω = 219.860 }2000.0
q = 5.18492 AU  i = 54.120

COMETS C/2008 W11, C/2009 W12, C/2008 X5, C/2008 X6 (SOHO)

Further to IAUC 9028, additional presumed comets have been found on SOHO website images. C/2008 W11 and C/2008 W12 were slightly diffuse Kreutz sungrazers (peak magnitudes ∼ 8 and 8.5, respectively). C/2008 X5 (Meyer group) and C/2008 X6 (Marsden group) were both small and stellar in appearance (mag ∼ 7.5).

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 W11 Nov. 29.921 16^{h}24.^{m}0 −23^{o}25′ C2 EB 2009-C57
C/2008 W12  29.963 16 24.4 −23 23 C2 RM 2009-E60
C/2008 X5 Dec. 7.022 16 56.8 −23 57 C2 RK 2008-Y06
C/2008 X6  7.504 16 53.5 −22 17 C2 RK 2008-Y06

COMET 216P/LINEAR Further to IAUCs 9005, 9010, and 9020, comet P/2009 D1 = P/2001

CV_{8} (cf. IAUC 9021) has been given the permanent number 216P. (See also MPC 65277.)

2009 March 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green